An electron accelerated from a rest position through a potential difference of 30.0V. Find the de Broglie wavelength of the electron after it has passed through the potential difference 


1
Expert's answer
2021-04-16T07:26:43-0400

From conservation of energy:


"PE=KE,\\\space\\
qV=\frac12 mv^2,\\\space\\
v=\sqrt{\frac{2qV}{m}}."

De Broglie wavelength of the electron:


"\lambda=\frac{h}{mv}=\frac{h}{\sqrt{2mqV}}=224\text{ pm}."